OpenStudy (anonymous):

A pizza place offers 4 different cheeses and 10 different toppings. In how many ways can a pizza be made with 1 cheese and 4 toppings?

OpenStudy (campbell_st):

well you will have Cheese 1st Topping 2nd Topp 3rd Top 4th topp Choice choice choice choice choice 4 x 10 x 9 x 8 x 7 = so if you choose a topping from any 10, there will only be 9 choices for the 2nd topping then only 8 toppings for the 3rd choice etc. multiply it out for the number of arrangements.

OpenStudy (anonymous):

850 840 830 820 these are the options.

OpenStudy (campbell_st):

ok... so its the above calculation divided by the number of arrangements of the 4 toppings. \[\frac{4 \times 10 \times 9\times 8 \times 7}{4 \times 3 \times 2\times 1}\]

OpenStudy (kropot72):

^ That means the order in which each selection of 4 toppings is made does not matter.

OpenStudy (anonymous):

so 840?

OpenStudy (campbell_st):

yep.. looks good